# Çukurova Üniversitesi Rektörlüğü 01250 Sarıçam, Adana, Turkey

### v1

@@
11institutetext: RWTH Aachen University, III. Physikalisches Institut A, Aachen, Germany22institutetext: Çukurova University
Çukurova Üniversitesi Rektörlüğü 01250 Sarıçam, Adana, Turkey
33institutetext: The University of Alabama,
500 University Blvd East, Tuscaloosa 35401 AL, USA
44institutetext: National Technical University of Athens
28 Oktovriou (Patision) 42, 10682 Athens, Greece
55institutetext: Baylor University,
Waco 76706, TX, USA
66institutetext: Institute of High Energy Physics, Chinese Academy of Sciences, 19B Yuruanlu, Shijingshan District, Beijing, China, 10004977institutetext: Tsinghua University,
Beijing, 100084, China
88institutetext: University of Bahrain,
P.O. Box 32038, Bahrain
99institutetext: The Lebanese University,
14 Badaro, Museum, Beirut, Lebanon
1010institutetext: Bogazici University,
Bebek 34342, Istanbul, Turkey
$1$institutetext: Boston University,
Boston, Massachusetts, USA
1212institutetext: University of Bristol,
Beacon House, Queens Road, Bristol BS8 1QU, UK
1313institutetext: Brown University,
182 Hope Street, Providence 02912, RI, USA
1414institutetext: Université Libre de Bruxelles,
Boulevard du Triomphe, B-1050 Bruxelles
1515institutetext: California Institute of Technology,
Pasadena, CA 91125, USA
1616institutetext: Carnegie Mellon University,
5000 Forbes Ave, Pittsburgh 15213, PA, USA
1717institutetext: CERN,
Espl. des Particules 1, 1211 Geneve 23, Switzerland
1818institutetext: National Central University,
Chung-Li, Taiwan, ROC
1919institutetext: Northern Illinois University,
1425 W. Lincoln Hwy., DeKalb 60115, IL, USA
2020institutetext: Deutsches Elektronen-Synchrotron DESY,
Notkestr. 85 22607, Hamburg, Germany
2121institutetext: Joint Institute for Nuclear Reaserch, JINR,
Joliot Curie 6, 141980 Dubna, Russia
$2$institutetext: University of Dundee,
Nethergate, Dundee, DD1 4HN, Scotland, UK
2323institutetext: Fermilab,
Wilson Road, Batavia 60510, IL, USA
2424institutetext: Florida State University,
600 W. College Ave., Tallahassee 32306, FL, USA
2525institutetext: University of Helsinki,
Gustaf Hällströminkatu 2, 00560 Helsinki, Finland
2626institutetext: The University of Hamburg, Institut für Experimentalphysik,
Luruper Chaussee 149, 22761 Hamburg, Germany
2727institutetext: Institut für Hochenergiephysik,
Nikolsdorfer Gasse 18, 1050 Wien, Vienna, Austria
2828institutetext: Imperial College,
Prince Consort Road SW7 2AZ, London, United Kingdom
2929institutetext: Indian Institute of Technology Madras,
60036 Chennai, India
3030institutetext: The University of Iowa,
203 Van Allen Hall, Iowa City, 52242, Iowa, USA
3131institutetext: National Centre for Physics,Quaid-I-Azam University,
Islamabad-44000, Pakistan.
3232institutetext: Istanbul Technical University,
Maslak, 80626 Istanbul, Türki̇ye.
$3$institutetext: Istanbul University,
TR-34452, Beyazit, Istanbul, Türki̇ye.
3434institutetext: Kansas State University,
116 Cardwell Hall, Manhattan, KS 66506, USA.
3535institutetext: Institut für Experimentelle Teilchenphysik,
Karlsruher Institut für Technologie, Wolfgang-Gaedestrasse 1, D-76131, Karlsruhe.
3636institutetext: King Abdullah University of Science and Technology, Thuwal 23955-6900, Saudi Arabia.3737institutetext: Institute for Scintillation Materials of National Academy of Science of Ukraine,
60 Lenina Ave, 61001 Kharkiv, Ukraine.
3838institutetext: NSC Kharkiv Institute of Physics and Technology,
1 Akademichna St., 61108 Kharkiv, Ukraine
3939institutetext: Saha Institute of Nuclear Physics,
HBNI, Bidhan Nagar, 700 064 Kolkata, India
4040institutetext: LIP,
Avenida Prof. Gama Pinto, n 2, 1649-003, Lisbon, Portugal
4141institutetext: National Centre for Particle Physics,
University of Malaya,Kuala Lumpur 50603,Malaysia
4242institutetext: The University of Maryland,
College Park 20742, MD, USA
4343institutetext: INFN, Università Degli Studi Milano-Bicocca,
Piazza della Scienza 3, I-20126 Milano
$4$institutetext: The University of Minnesota,
116 Church Street SE, Minneapolis 55405, MN, USA
4545institutetext: MIT, Laboratory for Nuclear Science, 77 Mass Ave, Cambridge, MA 02139, USA4646institutetext: Nanjing Normal University,
1 Wenyuan Rd., Qixia District, Nanjing, Jiangsu province, P.R. China 210023
4747institutetext: Northwestern University,
2145 Sheridan Rd, Evanston 60208, IL, USA
4848institutetext: University of Notre Dame,
Notre Dame 46556, IN, USA
4949institutetext: Laboratoire OMEGA CNRS/IN2P3,
Route de Saclay 91128, Ecole Polytechnique, France
5050institutetext: University of Montenegro,
Cetinjska br. 2, 81000 Podgorica, Crna Gora, Montenegro
5151institutetext: Laboratoire Leprince-Ringuet CNRS/IN2P3,
Route de Saclay, 91128 Ecole Polytechnique Cedex, France
5252institutetext: Indian Institute of Science Education and Research,
Dr. Homi Bhabha Road 411008, Pune, India
5353institutetext: Riga Technical University,
6A Kipsalas Street, Riga LV-1048, Latvia
5454institutetext: University of Rochester,
Campus Box 270171, Rochester 14627, NY, USA
$5$institutetext: CEA Paris-Saclay,
IRFU, Batiment 141,91191, Gif-Sur-Yvette Paris, France
5656institutetext: University of Split FESB ,
R. Boskovica 32, HR-21000, Split, Croatia
5757institutetext: Faculty of Science, University of Split,
R. Boskovica 33, 21000 Split, Croatia
5858institutetext: Bethel University,
3900 Bethel Drive, St. Paul, MN 55112, USA
5959institutetext: National Taiwan University,
10617, Taipei, Taiwan
6060institutetext: Texas Tech University,
Department of Physics and Astronomy, Lubbock 79409, TX, USA
6161institutetext: Georgian Technical University,
Kostava str, Tiiblisi, 77, 0160, Georgia.
6262institutetext: Tata Inst. of Fundamental Research-A,
Homi Bhabha Road, Mumbai 400005, India
6363institutetext: Tata Inst. of Fundamental Research-B,
Homi Bhabha Road, Mumbai 400005, India
6464institutetext: The University of California Santa Barbara,
Broida Hall, Santa Barbara 93106, CA, USA
6565institutetext: The University of Wisconsin,
Madison, WI, USA
$6$institutetext: Yildiz Technical University,
Esenler 34220, Istanbul, Turkey
6767institutetext: Zhejiang University,
866 Yuhangtang Rd, Hangzhou, Zhejiang, China
6868institutetext: Affiliated with an institute or an international laboratory covered by a cooperation agreement with CERN.
Using graph neural networks to reconstruct charged pion showers in the CMS High Granularity Calorimeter
A. Alpana 52    N. Kumar 42    M. Aamir 9    B. Acar 61    G. Adamov 24    T. Adams 18    C. Adloff 21    S. Afanasiev 68    S. Afanasiev 63    C. Agrawal 52    C. Agrawal 31    A. Ahmad 31    H. Ahmed 31    S. Akbar 60    N. Akchurin 66    B. Akgul 10    B. Akgun 10    R. Akpinar 33    E. Aktas 24    A. Al Kadhim 21    V. Alexakhin 68    V. Alexakhin 20    J. Alimena 16    J. Alison 36    W. Alshehri 17    P. Alvarez Dominguez 23    M. Alyari 17    C. Amendola 31    R. Amir 17    S. Andersen 68    Y. Andreev 17    P. Antoszczuk 10    U. Aras 35    L. Ardila 17    P. Aspell 19    M. Avila 9    I. Awad 2    Z. Azimi 17    A. Aznar Pretel 20    O. Bach 28    R. Bainbridge 23    A. Bakshi 3    B. Bam 65    S. Banerjee 17    D. Barney 51    F. Beaudette 51    F. Beaujean 51    E. Becheva 29    P. Behera 42    A. Belloni 27    T. Bergauer 55    M. Besancon 54    O. Bessidskaia Bylund 62    L. Bhatt 18    D. Bhowmil 20    F. Blekman 68    P. Blinov 28    P. Bloch 54    A. Bodek 68    a. Boger 51    A. Bonnemaison 55    F. Bouyjou 64    L. Brennan 17    E. Brondolin 35    A. Brusamolino 50    I. Bubanja 3    A. Buchot Perraguin 3    P. Bunin 56    A. Burazin Misura 64    A. Butler-nalin 33    Z. Caglar 32    A. Cakir 49    S. Callier 3    S. Campbell 17    K. Canderan 32    K. Cankocak 51    A. Cappati 18    S. Caregari 64    S. Carron 60    C. Carty 51    A. Cauchois 59    L. Ceard 66    S. Cerci 18    P. Chang 63    R. Chatterjee 27    S. Chatterjee 63    P. Chattopadhyay 4    T. Chatzistavrou 31    M. Chaudhary 62    A. Chauhan 59    J. Chen 18    J. Chen 42    Y. Chen 18    K. Cheng 23    H. Cheung 62    J. Chhikara 51    A. Chiron 51    M. Chiusi 61    D. Chokheli 3    R. Chudasama 12    E. Clement 17    S. Coco Mendez 56    D. Coko 32    K. Coskun 55    F. Couderc 44    B. Crossman 46    Z. Cui 51    T. Cuisset 23    G. Cummings 28    E. Curtis 45    M. D’Alfonso 28    J. D-hler-ball 37    O. Dadazhanova 60    J. Damgov 28    I. Das 39    S. Das Gupta 28    P. Dauncey 17    A. David Tinoco Mendes 28    G. Davies 51    O. Davignon 54    P. de Barbaro 49    C. De La Taille 20    M. De Silva 51    A. De Wit 30    P. Debbins 17    M. Defranchis 55    E. Delagnes 55    P. Devouge 62    C. Dewangan 23    G. Di Guglielmo 17    L. Diehl 30    K. Dilsiz 32    G. Dincer 5    J. Dittmann 27    M. Dragicevic 6    D. Du 21    B. Dubinchik 68    B. Dubinchik 62    S. Dugad 49    F. Dulucq 2    I. Dumanoglu 17    M. Dunser 33    B. Duran 39    S. Dutta 16    V. Dutta 19    A. Dychkant 42    T. Edberg 51    I. Ehle 49    A. El Berni 41    F. Elias 42    S. Eno 66    E. Erdogan 66    B. Erkmen 21    Y. Ershov 68    Y. Ershov 16    E. Ertorer 49    S. Extier 51    L. Eychenne 66    Y. Fedar 28    G. Fedi 17    J. Figueiredo De Sa Sousa De Almeida 51    B. Fontana Santos Alves 44    E. Frahm 19    K. Francis 23    J. Freeman 17    T. French 10    E. Gülmez 20    F. Gaede 23    P. Gandhi 55    S. Ganjour 54    A. Garcia-Bellido 51    F. Gastaldi 39    L. Gazi 23    Z. Gecse 17    H. Gerwig 55    O. Gevin 63    S. Ghosh 51    S. Ghosh 17    K. Gill 3    S. Gleyzer 56    N. Godinovic 10    M. Goek 20    P. Goettlicher 24    R. Goff 21    A. Golunov 68    A. Golunov 10    B. Gonultas 49    J. González Martinez 21    N. Gorbounov 68    N. Gorbounov 13    L. Gouskos 17    A. Gray 64    C. Grieco 25    S. Groenroos 17    D. Groner 17    S. Gronroos 17    A. Gruber 23    A. Grummer 55    F. Guilloux 2    Y. Guler 32    A. Gungordu 6    J. Guo 47    K. Guo 2    E. Gurpinar Guler 23    H. Gutti 66    A. Guvenli 33    S. Hale 68    Y. Halkin 60    G. Hamilton Ilha Machado 54    H. Hare 5    K. Hatakeyama 48    A. Heering 60    V. Hegde 13    U. Heintz 13    N. Hinton 20    A. Hinzmann 23    J. Hirschauer 15    D. Hitlin 6    B. Hou 6    X. Hou 28    A. Howard 42    C. Howe 59    H. Hsieh 59    T. Hsu 6    H. Hua 35    F. Hummer 33    H. Ilknur 31    M. Imran 64    J. Incandela 66    E. Iren 66    B. Isildak 22    P. Jackson 44    W. Jackson 63    S. Jain 29    P. Jana 17    J. Jaroslavceva 62    S. Jena 64    A. Jige 64    P. Jordano 23    U. Joshi 34    K. Kaadze 36    A. Kafizov 51    L. Kalipoliti 16    A. Kallil Tharayil 17    O. Kaluzinska 29    S. Kamble 68    A. Kaminskiy 16    M. Kanemura 9    H. Kanso 59    Y. Kao 50    A. Kapic 44    C. Kapsiak 21    V. Karjavine 68    V. Karjavine 59    S. Karmakar 68    A. Karneyeu 10    M. Kaya 33    B. Kaynak 2    A. Kays Topaksu 60    Y. Kazhykarim 14    F. Khan 68    A. Khudiakov 35    J. Kieseler 24    R. Kim 23    T. Klijnsma 16    E. Kloiber 35    M. Klute 17    Z. Kocak 62    K. Kodali 24    K. Koetz 24    T. Kolberg 66    O. Kolcu 29    J. Komaragiri 20    M. Komm 27    . Kopsalis 35    H. Krause 17    M. Krawczyk 62    T. Krishnaswamy Vinayakam 64    K. Kristiansen 56    A. Kristic 44    M. Krohn 42    B. Kronheim 20    K. Kruger 62    C. Kudtarkar 17    S. Kulis 55    M. Kumar 63    S. Kumar 25    R. Kumar Verma 60    S. Kunori 68    A. Kunts 18    C. Kuo 21    A. Kurenkov 68    A. Kurenkov 60    V. Kuryatkov 64    S. Kyre 13    J. Ladenson 60    K. Lamichhane 13    G. Landsberg 28    J. Langford 20    A. Laudrain 24    R. Laughlin 35    J. Lawhorn 51    O. Le Dortz 60    S. Lee 53    A. Lektauers 56    D. Lelas 17    M. Leon 38    L. Levchuk 64    A. Li 20    J. Li 59    Y. Li 7    Z. Liang 6    H. Liao 20    K. Lin 18    W. Lin 67    Z. Lin 23    D. Lincoln 17    L. Linssen 68    A. Litomin 51    G. Liu 6    Y. Liu 26    A. Lobanov 55    V. Lohezic 17    T. Loiseau 67    C. Lu 59    R. Lu 18    S. Lu 23    P. Lukens 47    M. Mackenzie 28    A. Magnan 51    F. Magniette 51    A. Mahjoub 44    D. Mahon 63    G. Majumder 68    V. Makarenko 21    A. Malakhov 68    A. Malakhov 17    L. Malgeri 28    S. Mallios 62    C. Mandloi 60    A. Mankel 17    M. Mannelli 44    J. Mans 23    C. Mantilla 24    G. Martinez 51    C. Massa 64    P. Masterson 17    M. Matthewman 21    V. Matveev 68    V. Matveev 62    S. Mayekar 68    I. Mazlov 17    A. Mehta 30    A. Mestvirishvili 47    Y. Miao 20    G. Milella 62    I. Mirza 62    P. Mitra 17    S. Moccia 62    G. Mohanty 17    F. Monti 17    F. Moortgat 16    S. Murthy 56    J. Music 48    Y. Musienko 42    S. Nabili 62    S. Nayak 58    J. Nelson 63    A. Nema 17    I. Neutelings 20    J. Niedziela 68    A. Nikitenko 23    D. Noonan 17    M. Noy 10    K. Nurdan 68    S. Obraztsov 51    C. Ochando 30    H. Ogul 24    J. Olsson 30    Y. Onel 33    S. Ozkorucuklu 59    E. Paganis 16    P. Palit 67    R. Pan 52    S. Pandey 17    F. Pantaleo 42    C. Papageorgakis 12    S. Paramesvaran 42    M. Paranjpe 63    S. Parolia 22    A. Parsons 54    P. Parygin 16    M. Paulini 45    C. Paus 24    K. Peñaló Castillo 23    K. Pedro 56    V. Pekic 60    T. Peltola 67    B. Peng 43    A. Perego 17    D. Perini 40    A. Petrilli 18    H. Pham 51    T. Pierre-Emile 63    S. Podem 38    V. Popov 51    L. Portales 55    L. Portales 33    O. Potok 28    P. Pradeep 63    R. Pramanik 24    H. Prosper 56    M. Prvan 17    S. Qasim 17    H. Qu 17    T. Quast 35    A. Quiroga Trivi?o 55    L. Rabour 50    N. Raicevic 62    H. Rajpoot 31    M. Rao 17    K. Rapacz 1    W. Redjeb 20    M. Reinecke 44    M. Revering 16    A. Roberts 11    J. Rohlf 17    P. Rosado 28    A. Rose 45    S. Rothman 18    P. Rout 17    M. Rovere 3    P. Rumerio 44    R. Rusack 20    L. Rygaard 17    V. Ryjov 37    S. Sadivnycha 55    M. Sahin 66    U. Sakarya 51    R. Salerno 44    R. Saradhy 63    M. Saraf 17    K. Sarbandi 10    M. Sarkisla 21    I. Satyshev 68    I. Satyshev 31    N. Saud 51    J. Sauvan 58    G. Schindler 1    A. Schmidt 30    I. Schmidt 47    M. Schmitt 56    A. Sculac 57    T. Sculac 68    A. Sedelnikov 28    C. Seez 20    F. Sefkow 20    D. Selivanova 17    M. Selvaggi 68    V. Sergeychik 31    M. Shahid 63    P. Sharma 6    R. Sharma 52    S. Sharma 62    M. Shelake 23    A. Shenai 18    C. Shih 63    R. Shinde 68    D. Shmygol 28    R. Shukla 17    E. Sicking 17    P. Silva 66    E. Simsek 62    B. Sirasva 51    Y. Sirois 6    S. Song 67    Y. Song 55    G. Soudais 42    S. Sriram 34    R. St Jacques 17    A. Stahl Leiton 17    A. Steen 16    J. Stein 23    J. Strait 44    N. Strobbe 59    X. Su 21    E. Sukhov 68    E. Sukhov 8    A. Suleiman 66    D. Sunar Cerci 62    P. Suryadevara 63    K. Swain 23    C. Syal 2    B. Tali 52    K. Tanay 18    W. Tang 31    A. Tanvir 6    J. Tao 51    A. Tarabini 10    T. Tatli 34    R. Taylor 66    Z. Taysi 23    G. Teafoe 59    C. Tee 16    W. Terrill 49    D. Thienpont 62    R. Thomas 55    M. Titov 22    C. Todd 24    E. Todd 35    M. Toms 17    J. Troska 59    L. Tsai 61    Z. Tsamalaidze 59    D. Tsionou 4    G. Tsipolitis 17    M. Tsirigoti 6    R. Tu 66    S. Tural Polat 60    S. Undleeb 3    E. Usai 2    E. Uslan 21    V. Ustinov 68    V. Ustinov 64    T. Vami 51    E. Vernazza 37    O. Viahin 24    O. Viazlo 17    P. Vichoudis 29    A. Vijay 28    T. Virdee 23    E. Voirin 28    M. Vojinovic 21    N. Voytishin 68    N. Voytishin 24    A. Wade 17    D. Walter 6    C. Wang 6    F. Wang 47    J. Wang 67    K. Wang 23    X. Wang 7    X. Wang 7    Y. Wang 6    Z. Wang 51    E. Wanlin 48    M. Wayne 30    J. Wetzel 23    A. Whitbeck 23    R. Wickwire 64    D. Wilmot 5    J. Wilson 59    H. Wu 67    M. Xiao 64    J. Yang 10    B. Yazici 67    Y. Ye 66    T. Yetkin 3    R. Yi 24    R. Yohay 6    T. Yu 6    C. Yuan 6    X. Yuan 66    O. Yuksel 20    I. YushmanoV 41    I. Yusuff 51    A. Zabi 22    D. Zareckis 21    A. Zarubin 68    A. Zarubin 17    P. Zehetner 51    A. Zghiche 6    C. Zhang 64    D. Zhang 6    H. Zhang 6    J. Zhang 24    J. Zhang 6    Z. Zhang 6    X. Zhao 67    J. Zhong 16    Y. Zhou km.alpana@cern.ch
@@ 1 Introduction
Among other physics cases, the high-luminosity running of the LHC (HL-LHC) will provide the data required to measure the Higgs couplings to a precision of approximately 1% and a first measurement of the its triple self-coupling bib.HLLHCYellowReport . Operations of the HL-LHC will start towards the end of this decade with a planned instantaneous luminosity of 5 $\times$ 1034 cm-2s-1 and an average number of interactions (pileup) per bunch crossing of 140 or more bib.cms-hgcalphase2-tdr . The endcap calorimeters of the CMS experiment that cover a pseudorapidity range of $3.01.5<|\eta|<$ 3.0 will be replaced with two high granularity calorimeters (HGCAL) designed to withstand the radiation exposure expected in the ten or more years of the HL-LHC operations bib.cms-phase2-techproposal . In the technical design report bib.cms-hgcalphase2-tdr , each HGCAL is a sampling calorimeter with a 28-layer electromagnetic section (CE-E) followed by a 22-layer hadronic section (CE-H). Recently the number of the CE-E and CE-H layers has been reduced to 26 and 21, respectively. The active detectors in the CE-E and CE-H are hexagonal 8-inch silicon sensors subdivided into 0.56 and 1.26 $\,\text{cm}^{2}$ hexagonal cells, to be used in the regions of higher and lower radiations levels, respectively. In the regions away from beam pipe and deep into the calorimeter, where radiation levels are sufficiently low, plastic scintillator cells ranging in size 5 to 25 $\,\text{cm}^{2}$, which are directly read out by silicon photomultipliers are used. In all, the total number of detector channels will be in excess of six million. This fine segmentation has been chosen for efficient particle identification, particle-flow reconstruction, and pileup rejection.
@@
In 2018 a prototype of the HGCAL calorimeter was exposed to the beam in the H2 beamline from the CERN SPS bib.cern-h2-manual . The prototype consisted of a 28-layer electromagnetic section with 6-inch silicon sensors subdivided into $1.1\,\text{cm}^{2}$ hexagonal detector channels and a 12-layer hadronic section, also equipped with silicon sensors. Digitization of the signal in each channel was performed with the SKIROC2-CMS ASIC bib.Skiroc2cms , which measured both amplitude and the time-of-arrival of the signal. The CALICE Analog Hadronic Calorimeter (AHCAL) bib.calice-ahcal-construction ; bib.calice-ahcal-2018tb-paper consisting of 39 layers of plastic scintillator tiles sandwiched between steel plates was placed at the back. Together they made a reasonable approximation of the final HGCAL design. To best approximate the sampling in the hadronic section of the HGCAL, only every fourth layer of the AHCAL was used in the analysis described here. The prototype was exposed to beams of positrons and pions with momenta between 20 and 300  GeV, and muons with momenta of 200  GeV. Details of the mechanical design or the readout electronics can be found in Refs. bib.cern-h1paper ; bib.cern-h2paper . The performance of the prototype calorimeter with positron and with pions beams based on the conventional weighted summation method has been reported previously in Refs. bib.hgcal-2018-positrons and bib.hgcal-2018-pions , and the results have been used to benchmark and to demonstrate the performance of a Geant4 simulation bib.GEANT4 of the prototype.
@@
When hadrons are incident on a calorimeter detector, there are a wide variety of nuclear processes that take place as the energy is absorbed. This leads to large event-by-event fluctuations in the types and multiplicity of shower particles, in the spatial distribution of their energy deposits, and in the fraction of invisible energy lost to nuclear binding energy, all of which depend on the energy of the incident particle. The energy reconstruction of these showers can be further complicated by energy leakage due to insufficient coverage of the detector or energy losses due to missing channels. Given the multidimensional nature of the problem of reconstructing the energy of hadronic showers, traditional algorithms such as the weighted summation method (WS method) described in bib.hgcal-2018-pions , are often complex, involve ad-hoc choices of parameterization, and do not fully mitigate the effects mentioned above. On the other hand, graph neural networks (GNNs) are particularly well suited to the application of reconstructing showers in HGCAL, since the number of cells active in hadronic showers are typically a few hundred out of some tens of thousand cells in the region and the spatial distributions of the active cells varies considerably event-to-event bib.gnns-particlereco .
@@
In this publication, we present a novel algorithm to reconstruct the energy of charged pion showers with a GNN, with a dynamic reduction network (DRN) architecture bib.drn-pointclouds . A significant improvement in the energy resolution compared to the WS method is achieved using this algorithm. The improvements to the performance are validated using both the Geant4 simulation and the beam test data collected in 2018. We also show how the algorithm performs for a calorimeter with a geometry similar to the final HGCAL detector.
@@
We describe briefly the experimental setup of the 2018 prototype beam tests in Section 2. The DRN model and procedure of energy reconstruction using the DRN is described in Section 3. Performance comparisons with the Monte Carlo and with the beam test data are presented in Section 4 and 5, respectively. The performance of the algorithm when applied to simulated events in the HGCAL geometry is summarized in Section 6.

### v3

# Using graph neural networks to reconstruct charged pion showers in the CMS High Granularity Calorimeter

@@ 1 Introduction
Among other physics cases, the high-luminosity running of the LHC (HL-LHC) will provide the data required to measure the Higgs couplings to a precision of approximately 1% and a first measurement of its triple self-coupling [1]. Operations of the HL-LHC will start towards the end of this decade with a planned instantaneous luminosity of 5 $\times$ 1034 cm-2s-1 and an average number of interactions (pileup) per bunch crossing of 140 or more [2]. The endcap calorimeters of the CMS experiment that cover a pseudorapidity range of $3.01.5<|\eta|<$ 3.0 will be replaced with two high granularity calorimeters (HGCAL) designed to withstand the radiation exposure expected in the ten or more years of the HL-LHC operations [3]. In the technical design report [2], each HGCAL is a sampling calorimeter with a 28-layer electromagnetic section (CE-E) followed by a 22-layer hadronic section (CE-H). Recently the number of the CE-E and CE-H layers has been reduced to 26 and 21, respectively. The active detectors in the CE-E and CE-H are hexagonal 8-inch silicon sensors subdivided into 0.56 and 1.26 $\,\text{cm}^{2}$ hexagonal cells, to be used in the regions of higher and lower radiation levels, respectively. In the regions away from beam pipe and deep into the calorimeter, where radiation levels are sufficiently low, plastic scintillator cells ranging in size 5 to 25 $\,\text{cm}^{2}$, which are directly read out by silicon photomultipliers are used. In all, the total number of detector channels will be in excess of six million. This fine segmentation has been chosen for efficient particle identification, particle-flow reconstruction, and pileup rejection.
@@
In 2018 a prototype of the HGCAL calorimeter was exposed to the beam in the H2 beamline from the CERN SPS [4]. The prototype consisted of a 28-layer electromagnetic section with 6-inch silicon sensors subdivided into $1.1\,\text{cm}^{2}$ hexagonal detector channels and a 12-layer hadronic section, also equipped with silicon sensors. Digitization of the signal in each channel was performed with the SKIROC2-CMS ASIC [5], which measured both amplitude and time-of-arrival of the signal. The CALICE Analog Hadronic Calorimeter (AHCAL) [6, 7] consisting of 39 layers of plastic scintillator tiles sandwiched between steel plates were placed at the back. Together they made a reasonable approximation of the final HGCAL design. To best approximate the sampling in the hadronic section of the HGCAL, only every fourth layer of the AHCAL was used in the analysis described here. The prototype was exposed to beams of positrons and pions with momenta between 20 and 300  GeV, and muons with momenta of 200  GeV. Details of the mechanical design or the readout electronics can be found in Refs. [8, 9]. The performance of the prototype calorimeter using positron and pion beams, evaluated with the conventional weighted summation method, has been previously reported in Refs. [10] and [11]. These results have been used to benchmark and to demonstrate the accuracy of the Geant4 simulation [12] of the prototype.
@@
When hadrons are incident on a calorimeter detector, there are a wide variety of nuclear processes that take place as the energy is absorbed. This leads to large event-by-event fluctuations in the types and multiplicity of shower particles, in the spatial distribution of their energy deposits, and in the fraction of invisible energy lost to nuclear binding energy, all of which depend on the energy of the incident particle. The energy reconstruction of these showers can be further complicated by energy leakage due to insufficient detector coverage or energy losses due to missing channels. Given the multidimensional nature of the problem of reconstructing the energy of hadronic showers, traditional algorithms such as the weighted summation method (WS method) described in [11], are often complex, involve ad-hoc choices of parameterization, and do not fully mitigate the effects mentioned above. On the other hand, graph neural networks (GNNs) are particularly well-suited for reconstructing showers in HGCAL, as the number of active cells in hadronic showers is typically a few hundred out of tens of thousands of cells in the region, and the spatial distributions of the active cells vary considerably from event to event [13].
@@
In this publication, we present a novel algorithm to reconstruct the energy of charged pion showers with a GNN based dynamic reduction network (DRN) architecture [14]. A significant improvement in the energy resolution compared to the WS method is achieved using this algorithm. The improvements to the performance are validated using both the Geant4 simulation and the beam test data collected in 2018. We also show how the algorithm performs for a calorimeter with a geometry similar to the final HGCAL detector.
@@
We describe briefly the experimental setup of the 2018 prototype beam tests in Section 2. The DRN model and procedure of energy reconstruction using the DRN are described in Section 3. Performance comparisons with the Monte Carlo and with the beam test data are presented in Section 4 and 5, respectively. The performance of the algorithm when applied to simulated events in the HGCAL geometry is summarized in Section 6.